OpenStudy (anonymous):

At what approximate rate (in cubic meters per minute) is the volume of a sphere changing at the instant when its surface area is 3 square meters and the radius is increasing at the rate of 1/5 meters per minute?

OpenStudy (turingtest):

V=(4/3)pi*r^3 A=4pi*r^2=3 and we are given that dr/dt=1/5 now compute dV/dt to find the rate change of volume and watch what formula you get (remember to use the chain rule because this is the derivative w/respect to t)

OpenStudy (anonymous):

am i finding the derivative for surface area or volume?

OpenStudy (turingtest):

volume with respect to time dV/dt is what the question asks for

OpenStudy (anonymous):

oh thats right.

OpenStudy (anonymous):

so would it be 4pi(r)^2dr/dt?

OpenStudy (turingtest):

the chain rule is going to give exactly now what does the first part of that formula look like?

OpenStudy (turingtest):

[4pi(r)^2]dr/dt ^^^^^^^ familiar?

OpenStudy (anonymous):

yes surface area formula!

OpenStudy (turingtest):

so we can just plug in A=3 for that and dr/dt=1/5 and we're done :)

OpenStudy (turingtest):

plus now you know that for a sphere A=dV/dr which is a nice thing to know

OpenStudy (anonymous):

3=4pi(r)^2(1/5)

OpenStudy (turingtest):

you put the 3 where dV/dt should be dV/dt=4pi(r)^2(dr/dt)=A(dr/dt) ...now plug in for A and dr/dt

OpenStudy (anonymous):

3/5

OpenStudy (turingtest):

yup

OpenStudy (anonymous):

oh wow so it would 3/5 cubic meters per minute

OpenStudy (anonymous):

thank you!

OpenStudy (turingtest):

anytime!

OpenStudy (anonymous):

i have another problem i need help with if you can help me im putting it up :)